Calculate the vapor pressure at 25 degrees C of an aqueous solution that is 5.50% NaCl by mass. The pure vapor pressure of water at this temperature is 23.78 torr.
Answer #1
we have

Nacl = 5.50 / (23+35.5) = 5.50 / 58.5 = 0.0940 moles
Water = (100-5.50) / 18 = 5.25 moles

therefore solution contains = 0.0940 + 5.25 = 5.344 moles

Fraction of water = 5.25/5.344 = 0.9824

therefore, vapor pressure = 0.9824 x 23.78 = 23.36 torr